Radio station KBOB broadcasts at a frequency of 85.7 MHz on your dial using a radio waves that travel at 3.00X10^8 m/s. . Since nearly all of the station's audience is due south of the transmitter, the managers of KBOB don't want to waste any energy broadcasting to the east and west. They decide to build two towers, transmitting at exactly the same frequency, aligned on an east-west axis. For engineering reasons, the two towers must be at least 10 meters apart. What's the shortest distance between the towers that will eliminate all broadcast power to the east and west? show all work please.
Answer #1

wavelength of the wave = speed of wave/frequency of wave = 3*10^{8}/85.7*10^{6}= 3.5006 m

the distance should be such that all the wave along the west & east axis have destructive resonance with each other.

Thus, distance between the 2 towers should be = (n+1/2)*wavelength

where n is an integer.

thus, (3+1/2)*3.5006 = 12.2521 m

n > 2.3566

this implies n = 3,

Thus distance between the 2 towers = (3+1/2)*3.5006 = 12.2521 m